OpenStudy (anonymous):

Use a calculator to find the mean and standard deviation of the data. Round to the nearest tenth. 946, 726, 956, 519, 104, 415, 428, 457, 614, 201, 772, 801

OpenStudy (tyteen4a03):

To find the mean, simply add them all together and divide the sum by the number of elements in this list. Now to find the standard deviation, first find the distance of each element from the mean. Subtract the mean from each element and then square them all, and there goes your distance. Then, we'll need to find the variance. Add these distances up and then divide the sum by the number of elements in the list. Now take the square root of the variance and you have your SD. A one-go formula is \(\sigma = \sqrt{\frac{1}{N}\displaystyle\sum\limits_{i=1}^N (x_{i} - \mu)^2 }\), where \(x_{i}\) stands for each element, \(\mu\) stands for the mean and N stands for the number of elements in this list. It does exactly what the second paragraph said - only with a tidier (read: much more complicated to calculate) formula.